A 42.477 km child swings on a playground swing. If the swing hangs down 2.814 meters from the top bar, and the child swings out 20.88 degrees, what is the change in potential energy of the child?
Answer #1

At 20.88º the swing seat is a vertical distance L below the top bar
Cos20.88 = L/2.814 =====> L = 2.814*cos20.88 =====> L = 2.63 m
At 20.88º the seat is (2.814 - L) m vertically above it's rest position = 0.184 m

The child's height above ground = 0.184 m
Gain in PE = mg ∆h = 42.477  x 9.8 x 0.184 ======> ∆PE = 76.6 J